gunther invested $10,000 in two mutual funds. one of the funds rose 6% in one year , and the other rose9% in one year. if gunther's onvestment rose a total of $684 in one year , how much did he invest in each mutual fund?
Answer by addingup(3677)

How much at 6%? Let's call the amount x
How much at 9%? Well, 10,000-x because x is at 6%
0.06x+0.09(10,000-x) = 684
0.06x+900-0.09x = 684
-0.03 = -216
Divide -216/-0.03 to get your answer, and remember -/-= +